OpenStudy (anonymous):

Rewrite the rational exponent as a radical. 5 to the 3 over 4 power, to the 2 over 3 power

OpenStudy (jdoe0001):

\(\large \left( 5^{\frac{3}{4}} \right)^{\frac{2}{3}} ?\)

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

anyhow, keep in mind that \(\bf (a^n)^m\implies a^{n\cdot m}\qquad and\qquad a^{\frac{{\color{blue} n}}{{\color{red} m}}} \implies \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}\implies a^{\frac{{\color{blue} n}}{{\color{red} m}}}\)

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

thus \(\large \bf \left( 5^{\frac{3}{4}} \right)^{\frac{2}{3}}\implies 5^{\frac{3}{4}\cdot \frac{2}{3}}\implies 5^{\frac{{\color{blue}{ 3\cdot 2 }}}{{\color{red}{4\cdot 3 }}}}\)

OpenStudy (anonymous):

so after you multiply the exponents do you put it back in that form at the top^